A loaded Airbus A380-800 aircraft with a mass of 520 tons flies in the air. The wing area of ​​the aircraft is 841 M². Determine the mean pressure difference between the lower and upper sides of the aircraft wings.


1
Expert's answer
2021-03-09T15:28:26-0500
"P=\frac{F}{A}\\P=\frac{mg}{A}\\P=\frac{(9.8)(520000)}{841}=6060\ Pa"